Hello, I have a probability question and I don't know how to answer it if you could please show me I have several with your help I can solve them myself. Thank you so much.

The odds in favor of Sprinting Sally winning the horse race are 19:2. Determine the probability that Sprinting Sally wins the horse race
Thank you again

Answer by stanbon(75887)

The odds in favor of Sprinting Sally winning the horse race are 19:2. Determine the probability that Sprinting Sally wins the horse race
---
If "odds for an event (e)" are a:b
then P(e) = a/(a+b) P(not e) = b/(a+b)
-------------------------
Your problem:
P(Sally wins) = 19/(19+2) = 19/21
---------------------------------------
Cheers,
Stan H.